### v1

@@ Abstract
Geometric shape classification of vector polygons remains a non-trivial learning task in spatial analysis. Previous studies mainly focus on devising deep learning approaches for representation learning of rasterized vector polygons, whereas the study of discrete representations of polygons and subsequent deep learning approaches have not been fully investigated. In this study, we investigate a graph representation of vector polygons and propose a novel graph message-passing neural network (PolyMP) to learn the geometric-invariant features for shape classification of polygons. Through extensive experiments, we show that the graph representation of polygons combined with a permutation-invariant graph message-passing neural network achieves highly robust performances on benchmark datasets (i.e., synthetic glyph and real-world building footprint datasets) as compared to baseline methods. We demonstrate that the proposed graph-based PolyMP network enables the learning of expressive geometric features invariant to geometric transformations of polygons (i.e., translation, rotation, scaling and shearing) and is robust to trivial vertex removals of polygons. We further show the strong generalizability of PolyMP, which enables generalizing the learned geometric features from the synthetic glyph polygons to the real-world building footprints.
@@ 1 Introduction
Spatial objects are often conveniently represented as vector polygons, a data representation thus far neglected in deep learning research. Learning geometric-invariant features from vector polygons has the following requirements: (1) a generic data representation that encodes geometric features of polygons without information loss; and (2) a learning model built on this input data representation that enables learning latent geometric-invariant features robust to geometric transformations. Existing geospatial applications (i.e., shape coding and retrieval [6], building patter recognition [2, 1] and building grouping [7]) utilizing polygonal geometries motivate us to systemically study geometry encoding methods in conjunction with appropriate and robust learning architectures that enable learning transformation invariant features of spatial polygon geometries.
Veer et al. [2] proposed VeerCNN, a deep convolution model to learn convolutional features on fixed-size 1D vertex sequences of polygon vertices for building attribute recognition. The architectural limitations of CNNs attributable to shared-weights convolutional kernels followed by non-linear activation and fixed-size pooling layers (e.g, $mean$, $sum$ or $max$) only enable learning intermediate hierarchical features invariant to translation, with poor handling of rotation, scaling and shearing. Mai et al. [8] noted that CNN models learning on 1D sequences are also sensitive to (1) permutations of the feed-in order of polygon vertices (i.e., sensitive to loop origin invariance); and (2) the impact of trivial vertices on the exterior and interiors of polygons, defined by Mai et al. [8] as “[vertices] …where the addition or removal of the vertex have no effect on the geometric shape and topological properties of the outlines of the polygons”. In this study, we treat the addition or removal of trivial vertices of polygons as changes that do not alter the semantic information of the shape (captured by the human-assigned label), thus closely reflecting the relationship with the Gestalt principles.
@@
Graph convolutional autoencoders (GCAE) [6] extend graph convolution neural networks [16] by learning spectral convolution embeddings from polygon graphs and demonstrate the effectiveness of latent embedding. Bronstein et al. [13] discussed that the graph convolution layers aggregate node features from neighboring nodes with constant weights, largely limited by the topological structure of the input graphs. In vector polygons vertices on the linear rings have constant neighbours (i.e., left and right neighbour vertices on linear rings), leading to a reduction in the expressivity of graph convolution features learned from polygon graphs. In this study, we propose to leverage graph message-passing mechanisms [17] in graph-based learning models to learn highly expressive and robust latent features of polygons.
Our hypothesis is that graph representations of polygons enable the learning of robust latent features of polygons invariant to geometric transformations (i.e., rotation, scaling and shearing) when combined with graph message-passing models, thus supporting generalizability on data with domain shifts.
If the hypothesis holds, we suggest that the downstream tasks performed on the learned features of vector polygons lead to robust performance on shapes with varying amounts of trivial vertices. We present the results of a series of experiments designed to evaluate model robustness on polygons altered by geometric transformations, and investigate the generalizability of the findings on combinations of polygon representations and model architectures (permutation-invariant vs. translation-invariant models) for the classification of vector polygons with and without holes (inner linear rings).
After a literature review of related literature (Section 2) we present a synthetic dataset of polygonal shapes of high shape variability with known, robust human labels, based on latin alphabet character glyphs (Section 3.1.1). Despite a broad range of shape variations between fonts (in particular with respect to trivial vertices), characters must maintain strong recognisability attributable to the shapes’ Gestalt. We then experimentally benchmark the performance of learning models operating on three distinct data representations (1D sequence, set and graph) and varying amounts of geometric transformations (i.e., rotation, scaling and shearing). Our results (Section 4) document the models’ robustness to geometric transformations by benchmarking their performance on the synthetic glyph dataset. Building on these results, we further evaluate the generalizability of the models on a real-world building footprint dataset collected from the Open Street Map (OSM) [6].
@@
    a thorough investigation of the impact of distinct combinations of graph representation of polygons and message-passing learning architectures for the learning of geometric features robust to geometric transformations of shapes and removals of trivial vertices, grounded in a theoretically grounded set of desiderata;
@@
    the contribution of an open synthetic dataset of vector polygon shapes based on glyphs, enabling the benchmarking of geometric learning models; and
@@
    the experimental demonstration that graph representations of vector polygons in conjunction with message-passing neural networks achieve the most robust performance and generalization to real-world shape datasets.
@@ 2.1 Machine Learning with Vector Geometries
In a straightforward extension of computer vision approaches devised for image classification, current research on polygonal shape classification typically learns salient geometric features of shapes based on rasterized 2D vector geometries and traditional deep CNN architectures [18, 19].
The geospatial community has adopted these approaches for a variety of remote sensing tasks, including vector shape generation. The Microsoft open building footprints dataset [20] was generated by segmenting and subsequently vectorizing building polygons based on satellite and aerial imagery using deep semantic segmentation networks [21].
Such expertise in image classification naturally influenced approaches to learning tasks on vector shapes. Xu et al. [22] trained a deep convolution autoencoder to assess the quality of geographic information of rasterized building footprints collected from OpenStreetMap (OSM) [23]. Veer et al. [2] evaluated the classification performance of deep neural networks on spatial vector geometries with geometry vertices encoded as 1D sequences directly feed into deep learning models. While this shows the ability of CNNs and RNNs to achieve performances comparable to machine learning methods with hand-crafted features, these approaches neglect the desirable property of learning transformation invariant features of shapes.
Beyond the geospatial community, the typography and computer graphics communities also significantly handle vector shapes. Lopes et al. [24] developed a generative model with convolution layers as encoders, to generate vector text glyphs. Yet, this model is optimized to learn only scale-invariant features of font styles. Mino and Spanakis [25] and Sage et al. [26] generated vector icons and logos from rasterized images using deep generative autoencoders proposed by Radford et al. [27] and Odena et al. [28]. Furthermore, Ha and Eck [29] proposed a recurrent neural network (RNN), sketch-rnn, to generate stroke-based vector drawings of common objects, while later Carlier et al. [30] introduced a hierarchical generative network (DeepSVG) which uses Transformer blocks [31] as the encoder and decoder of a backbone architecture for scalable vector graphics (SVG) generation.
@@ 2.2 Vector Geometry as Discrete Data Representation
Recently, the progress of graph representation learning [32, 33, 16, 34], enabled studies re-formalized learning tasks of vector polygons as graph or set representation learning problems. Previous work by Yan et al. [1] investigated the building pattern classification problem utilizing graph convolution networks. They processed buildings into building clusters, based on geometric patterns of individual buildings. They discovered the relational representations between buildings in graphs built using clusters based on Delaunay triangulation and Minimun Spanning Tree. They then trained a graph convolution model [16] for a binary classification of regular and irregular building clusters. Similarly, Bei et al. [15] devised a spatial adaptive model that makes use of graph encoding, but applied to entire building clusters (i.e., buildings as graph nodes) and applied a graph convolution neural network [16] for group pattern recognition (i.e., inner-state, edge-state and free state building patterns).
Most recently, Yan et al. [6] presented a graph convolutional autoencoder, based on the graph convolution model [16], to analyze the shape of building polygons. Different from previous works, they proposed to represent building polygons as graphs with the vertices of boundary geometries of buildings as graph nodes and their neighbour vertices as description features. However, this method did not consider the relational structures between connected vertices in graphs. Hence, their experiments on building shape coding and empirical results of building shape similarity measurement suggest that the proposed method is sensitive to rotation of building polygons.
Liu et al. [35] proposed a deep point convolutional network (DPCN), a modification of a DGCNN [36], to recognize the geometric shapes of buildings in map space, with the convolution operator TriangleConv. Their method operates on the coordinates of vertices of the exterior boundaries of building polygons and the convolution operator TriangleConv extracts local geometric features from triangle features, where triangle features are given by $f_{\verb+TriangleConv+}(x_{i-1},x_{i},x_{i+1})=\{(x_{i}-x_{i-1}),(x_{i}-x_{i+1}),(x_{i-1}-x_{i+1})\}$, and $(x_{i-1},x_{i},x_{i+1})$ are adjacent coordinate features of a building polygon. This method, trained and evaluated on a building footprint dataset [6], achieves fair classification performances (i.e., accuracy) on geometric shapes of buildings. However, their method requires a pre-definition of local triangles constructed by the graph nodes and adjacency points, and attempts to aggregate the local triangle features (i.e., angles and areas) for updating node features. We argue that learning additional local triangle features of polygons repeatedly is redundant and increase computing overhead in model training, as pointed out by Tailor et al. [37]. Further, they do not evaluate the robustness and generalizability of permutation-invariant data representations and learning architectures to geometric transformations of spatial geometries.
@@
Mai et al. [8] recently proposed a general-purpose polygon encoding model for single and multi-part polygonal geometries. Their polygon encoding method, ResNet1D is a 1D CNN-based polygon encoder that captures local geometric features of spatial polygons and is deemed suitable for shape classification. ResNet1D encodes a polygon coordinate vertex $i$ as relative offset vectors from $k$ neighbouring vertices: $\{x_{i},x_{i-1}-x_{i},x_{i+1}-x_{i},\cdots,x_{i+k}-x_{i},x_{i-k}-x_{i}\}$. They contrast this encoding with a Non-Uniform Fourier Transformation (NUFT) method NUFTspec, based on the shape spectral domain. The NUFTspec aggregates global features of polygons as a NUFT representation and is a better encoding method for geometry topological relation predictions. Their work is thus parallel to our study. Here, we focus on the ability to support specific transformation invariant properties of combinations of learning architectures (permutation-variant CNN, set-based, graph-based and attention-based permutation-invariant models) and data representations (1D grid, set, graph and complete graph).
@@
While some of the above studies leverage graph representations and local geometric encoding of vector polygons and state-of-art graph representation learning models to solve tasks on vector shapes, they do not yet provide a grounded, principled justification for graph-based learning methods on vector polygons. We attempt to fill this gap here by providing experimental results grounded in a sound theoretical motivation, and demonstrate why discrete data representations in combination with permutation-invariant learning models provide desired robustness to vector shape learning tasks.
2.3 Graph Representation of Polygon
Polygons are point sets connected by lines, constituting a collection of clock-wise or counter clock-wise linear rings [38]. According to the Gestalt principles of invariance, the compact embedding of linear rings should ideally be invariant to geometric transformations (rotation, scale and shear). We define a graph $\mathcal{G}=(X,E)$, where $X$ is a node matrix that contains the point coordinates of geometry vertices and $E$ is an edge matrix that captures the connectivity of these vertices along the linear rings (boundaries of the outside and of any holes of polygons), hence encoding the node connectivity of the graph. The binary adjacency value $a_{ij}=1\in A$ is used to indicate the connectivity between nodes $i$ and $j$, with $A$ being a square adjacency matrix.
@@ 2.4 Graph Representation Learning
Considering a permutation matrix $\mathcal{P}$ operating on graph $\mathcal{G}$, the permutation matrix $\mathcal{P}$ alters the order of node feature matrix $X$ in $\mathcal{G}$ along with the adjacency matrix $A$, producing a perturbed graph $\mathcal{G}^{\prime}=(PX,PAP^{\intercal})=(X^{\prime},A^{\prime})$. The ideal permutation-invariant function $\mathcal{F}$ should satisfy $\mathcal{F}(X^{\prime},A^{\prime};\theta)=\mathcal{F}(X,A;\theta)$. We relate the idea of permutation matrix $\mathcal{P}$ to the aforementioned geometric transformations.
@@
The local differentiable function $\phi$ linearly transforms each node feature individually $x$ to latent space and the aggregation function sums over latent node features  [^1] The basic aggregation in this setting can also be averaging or max pooling., and $\rho$ is a global differentiable function applied on the summed node features followed by a non-linear activation function. In this simple setting, the outputs of $\mathcal{F}$ are invariant to the permutation of node features since the aggregation sum returns the same outputs for any input permutation: $Aggr_{sum}(x_{1},x_{2},....,x_{n})=Aggr_{sum}(x_{n},...,x_{2},x_{1})$.
This learning setting [9] makes a simple assumption on the permutations acting on input data, which can affect only individual node features without taking into account the connectivity within data (i.e., boundaries of polygon geometries). Graph convolution neural networks (GCN) by Kipf and Welling [16] learn convolutional features of node matrix $X$ with the normalized graph Laplacian matrix $\hat{A}=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}$, where $\tilde{A}=A+I$ is the adjacency matrix with a identity matrix (i.e., self-connections of nodes), and $\tilde{D}$ is the diagonal degree matrix. Intuitively, the graph Laplacian matrix describes the divergence of energy flowing from one place (source nodes) to another (target nodes) in a graph. Linking this back to the linear rings of polygons and graph representation, the graph Laplacian conceptually shows similar energy divergence for nodes on the same rings(i.e., polygons with holes) and for nodes that are close to each other (i.e, polygons without holes).
@@
2.5 Message-passing Neural Network
The principle of the message-passing mechanism is the aggregation of neighbour node features in graphs based on the “message” computed between the source and target nodes. We define the message-passing function operating on a single node as:

### v2

@@ Abstract
Geometric shape classification of vector polygons remains a challenging task in spatial analysis. Previous studies have primarily focused on deep learning approaches for rasterized vector polygons, while the study of discrete polygon representations and corresponding learning methods remains underexplored. In this study, we investigate a graph-based representation of vector polygons and propose a simple graph message-passing framework, PolyMP, along with its densely self-connected variant, PolyMP-DSC, to learn more expressive and robust latent representations of polygons. This framework hierarchically captures self-looped graph information and learns geometric-invariant features for polygon shape classification. Through extensive experiments, we demonstrate that combining a permutation-invariant graph message-passing neural network with a densely self-connected mechanism achieves robust performance on benchmark datasets, including synthetic glyphs and real-world building footprints, outperforming several baseline methods. Our findings indicate that PolyMP and PolyMP-DSC effectively capture expressive geometric features that remain invariant under common transformations, such as translation, rotation, scaling, and shearing, while also being robust to trivial vertex removals. Furthermore, we highlight the strong generalization ability of the proposed approach, enabling the transfer of learned geometric features from synthetic glyph polygons to real-world building footprints.
@@ 1 Introduction
Spatial objects are often conveniently represented as vector polygons, a discrete data representation thus far neglected in deep learning research. Learning geometric-invariant features from vector polygons has the following requirements: (1) a generic data representation that encodes geometric features of polygons without information loss; and (2) a learning model built on this input data representation that enables learning latent geometric-invariant features robust to geometric transformations. Existing geospatial applications (i.e., shape coding and retrieval [6], building patter recognition [2, 1] and building grouping [7]) utilizing polygonal geometries motivate us to systemically study geometry encoding methods in conjunction with appropriate and robust learning architectures that enable learning transformation invariant features of spatial polygon geometries.
Veer et al. [2] proposed VeerCNN, a deep convolution model to learn convolutional features on fixed-size 1D vertex sequences of polygon vertices for building attribute recognition. The architectural limitations of CNNs attributable to shared-weights convolutional kernels followed by non-linear activation and fixed-size pooling layers (e.g, $mean$, $sum$ or $max$) only enable learning intermediate hierarchical features invariant to translation, with poor handling of rotation, scaling and shearing. Mai et al. [8] noted that CNN models learning on 1D sequences are also sensitive to (1) permutations of the feed-in order of polygon vertices (i.e., sensitive to loop origin invariance); and (2) the impact of trivial vertices on the exterior and interiors of polygons, defined by Mai et al. [8] as “[vertices] …where the addition or removal of the vertex have no effect on the geometric shape and topological properties of the outlines of the polygons”. The addition or removal of trivial vertices in polygons does not alter the semantic information of the shape (as captured by the human-assigned label), closely aligning with the Gestalt principles.
@@
Graph convolutional autoencoders (GCAE) [6] extend graph convolutional neural networks (GCNs) [16] by learning spectral graph embeddings from polygon graphs, demonstrating the effectiveness of graph latent embeddings in polygon shape retrieval. Bronstein et al. [13] noted that spectral graph convolutions aggregate node features from neighboring nodes with constant weights, making them highly dependent on the topological structure of the input graphs. In vector polygons, vertices on linear rings have fixed neighbors (i.e., the left and right adjacent vertices), which reduces the expressivity of graph convolutional features learned from polygon graphs. In this study, we leverage graph message-passing mechanisms [17] and propose a simple graph message-passing network, PolyMP, along with its densely self-connected variant, PolyMP-DSC, to learn more expressive and robust latent representations of polygons. We hypothesize that combining graph-based representations of polygons with graph message-passing models facilitates learning robust latent features that remain invariant to geometric transformations (i.e., rotation, scaling, and shearing), thereby improving the generalizability of shape classification across datasets.
If our hypothesis holds, we suggest that downstream tasks performed on the learned features of vector polygons will exhibit robust performance on shapes with varying amounts of trivial vertices. We present a series of experiments designed to evaluate model robustness on polygons subjected to geometric transformations and investigate the generalizability of our findings across different combinations of polygon representations and model architectures (permutation-invariant vs. translation-invariant models) for classifying vector polygons with and without holes (inner linear rings).
Following a literature review (Section 2), we introduce a synthetic dataset of polygonal shapes with high shape variability and well-defined human labels, derived from Latin alphabet character glyphs (Section 4.1). Despite significant shape variations across fonts, particularly in terms of trivial vertices, characters maintain strong recognizability due to their Gestalt properties. We then benchmark the performance of learning models on three distinct data representations (1D sequence, set, and graph) under varying geometric transformations (i.e., rotation, scaling, and shearing). Our results (Section 5) document the models’ robustness to geometric transformations through performance evaluation on the synthetic glyph dataset. Building on these findings, we further assess model generalizability using a real-world building footprint dataset from OpenStreetMap (OSM) [6].
@@
    we provide a systematic investigation of vector polygon representations for geometric shape classification, highlighting the strengths of graph-based encoding over sequences and point sets;
@@
    the introduction of PolyMP and PolyMP-DSC, two lightweight graph message-passing architectures designed to learn geometric-invariant and robust latent representations;
@@
    the release of a synthetic dataset of vector polygons derived from character glyphs, enabling controlled benchmarking of geometric learning models; and
  4. 4.
    extensive experiments which show that the proposed message-passing models, PolyMP and PolyMP-DSC, significantly improve robustness to transformations and generalize well to real-world geospatial data.
Together, these contributions aim to provide a principled and practical foundation for learning geometric-invariant representations of spatial polygons—pushing forward the capabilities of shape-based analysis in geospatial domains.
@@ 2.1 Machine Learning with Vector Geometries
Building on computer vision approaches for image classification, polygonal shape classification has traditionally relied on learning geometric features from rasterized 2D vector geometries using deep CNN architectures [18, 19]. The geospatial community has adopted these methods for various remote sensing tasks, including vector shape generation. For example, the Microsoft Open Building Footprints dataset [20] was generated by segmenting and vectorizing building polygons from satellite and aerial imagery using deep semantic segmentation networks [21].
This expertise in image-based learning naturally influenced vector shape learning. Xu et al. [22] trained a deep convolutional autoencoder to assess the quality of rasterized building footprints collected from OpenStreetMap (OSM) [23]. Similarly, Veer et al. [2] evaluated deep neural networks for classifying spatial vector geometries, encoding geometry vertices as 1D sequences directly fed into deep models. While CNNs and RNNs performed comparably to hand-crafted feature-based methods, they lacked the ability to learn transformation-invariant shape features.
Beyond geospatial applications, typography and computer graphics research also leverage vector shape learning. Lopes et al. [24] developed a convolution-based generative model to create vector text glyphs, but it was optimized primarily for scale-invariant font style learning. Mino and Spanakis [25] and Sage et al. [26] generated vector icons and logos from rasterized images using deep generative autoencoders [27, 28]. Additionally, Ha and Eck [29] introduced Sketch-RNN, a recurrent neural network for generating stroke-based vector drawings, while Carlier et al. [30] proposed DeepSVG, a hierarchical generative model using Transformer-based encoding for scalable vector graphics (SVG).
@@ 2.2 Vector Geometry as Discrete Data Representation
Recent advances in graph representation learning [31, 32, 16, 33] have reformulated vector polygon learning as a graph-based or set-based problem. Yan et al. [1] applied graph convolution networks (GCNs) to building pattern classification, structuring buildings into clusters based on geometric relationships using Delaunay triangulation and Minimum Spanning Trees. Their GCN model classified building clusters into regular or irregular patterns. Similarly, Bei et al. [15] introduced a spatially adaptive model using graph encoding for group pattern recognition, where buildings served as nodes in a graph convolutional network (GCN).
Yan et al. [6] proposed a graph convolutional autoencoder (GCAE) for building shape analysis, encoding building polygons as graphs where boundary vertices acted as nodes. However, their method was sensitive to rotations, limiting its robustness for polygon shape retrieval. Addressing this limitation, Huang et al. [34] introduced a contrastive graph autoencoder to enable robust polygon shape matching and retrieval in large-scale vector datasets.
Liu et al. [35] developed a deep point convolutional network (DPCN), a modification of a DGCNN [36], for building shape recognition in map space. Their model introduced the TriangleConv operator, which extracts local geometric features from triangle-based representations of polygon vertices: $f_{\texttt{TriangleConv}}(x_{i-1},x_{i},x_{i+1})=\{(x_{i}-x_{i-1}),(x_{i}-x_{i+1}),(x_{i-1}-x_{i+1})\}$, where $(x_{i-1},x_{i},x_{i+1})$ are adjacent point features of a building polygon. Despite achieving competitive classification performance on building footprints [6], this method requires pre-defining local triangles of polygons and repeatedly computing redundant local geometric features (i.e., angles and areas) of polygons. Furthermore, it does not evaluate the robustness of learned representations to geometric transformations such as rotation, scaling, and vertex perturbations.
@@
Mai et al. [8] introduced ResNet1D, a 1D CNN-based polygon encoder that captures local geometric features for shape classification. It encodes polygon vertices using relative offsets from neighboring vertices: $\{x_{i},x_{i-1}-x_{i},x_{i+1}-x_{i},\cdots,x_{i+k}-x_{i},x_{i-k}-x_{i}\}$. They compared this encoding with NUFTSpec, a Non-Uniform Fourier Transform (NUFT)-based method that aggregates global shape features in the spectral domain. While ResNet1D focuses on shape classification, NUFTSpec is more suited for predicting topological relations. Their work aligns with our study in exploring different polygon encoding methods, but our focus is on evaluating transformation-invariant learning architectures across various representations (CNN, set-based, graph-based, and attention-based models).
@@
While previous studies explore graph-based and local geometric encoding for vector polygons, they often lack a principled justification for applying graph learning methods to polygon representations. We address this gap by providing: 1. Experimental validation grounded in theoretical motivation for robust polygon shape learning; 2. A principled framework demonstrating how permutation-invariant learning architectures combined with discrete data representations improve robustness to geometric transformations and vertex perturbations.
@@ 2.3 Graph Representation Learning
Given a permutation matrix $\mathcal{P}$ acting on a graph $\mathcal{G}$, it reorders the node feature matrix $X$ and the adjacency matrix $A$, producing a permuted graph representation $\mathcal{G}^{\prime}=(PX,PAP^{\intercal})=(X^{\prime},A^{\prime})$. The ideal permutation-invariant function $\mathcal{F}$ should satisfy $\mathcal{F}(X^{\prime},A^{\prime};\theta)=\mathcal{F}(X,A;\theta)$. We relate the idea of permutation matrix $\mathcal{P}$ to the aforementioned geometric transformations.
@@
The local differentiable function $\phi$ linearly transforms each node feature individually $x$ to latent space and the aggregation function sums over latent node features  [^1] The basic aggregation in this setting can also be averaging or max pooling., and $\rho$ is a global differentiable function applied on the summed node features followed by a non-linear activation function. In this simple setting, the outputs of $\mathcal{F}$ are invariant to the permutation of node features since the summation-based aggregation returns the same outputs for any input permutation: $Aggr_{sum}(x_{1},x_{2},....,x_{n})=Aggr_{sum}(x_{n},...,x_{2},x_{1})$.
This framework (i.e., DeepSet [9]) assumes that permutations affect only individual node features while disregarding their connectivity, which is crucial for polygonal structures. Graph convolution neural networks (GCN) by Kipf and Welling [16] learn convolutional features of node matrix $X$ with the normalized graph Laplacian matrix $\hat{A}=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}$, where $\tilde{A}=A+I$ is the adjacency matrix with a identity matrix (i.e., self-connections of nodes), and $\tilde{D}$ is the diagonal degree matrix. The graph Laplacian matrix describes the divergence of energy flowing from source nodes to target nodes, analogous to the spatial relationships in polygonal structures. Specifically, it captures connectivity patterns both within polygons (e.g., rings with holes) and between adjacent polygons.
@@
3 Approach
3.1 Graph Representation of Polygon
Polygons are sets of points connected by lines, forming a collection of clockwise or counterclockwise linear rings [38]. According to the Gestalt principles of invariance, the compact embedding of linear rings should ideally be invariant to geometric transformations such as rotation, scaling, and shear. We define a graph $\mathcal{G}=(X,E)$, where $X$ is a node matrix containing the coordinates of the geometry vertices, and $E$ is an edge matrix capturing the connectivity of these vertices along the linear rings (i.e., the boundaries of the polygon and any holes). The adjacency matrix $A$ encodes the connectivity between nodes, where a binary value $a_{ij}=1$ indicates that nodes $i$ and $j$ are connected.
3.2 Message-passing Encoder
The principle of message-passing involves aggregating neighboring node features in graphs based on messages computed between source and target nodes. We define the message-passing function operating on a single node as: